OpenStudy (anonymous):

solve by elimination 2x-y=19 3y=2x-29

OpenStudy (anonymous):

2x-y=19 -2x+3y=-29(signs reversed) ------------------ 2y = -10(2x and -2x cancels out) y=-5.. subsitiute y=-5 into first equation 2x-(-5)=19 2x+5=19 2x=14 x=7

OpenStudy (anonymous):

can you please explain it to me?

OpenStudy (anonymous):

which part did u not understand

OpenStudy (anonymous):

did you add them together or multiply?

OpenStudy (anonymous):

there are two equations. i subtracted the second equation. see the signs are reversed.

OpenStudy (anonymous):

that way i eliminated 2x

OpenStudy (anonymous):

ok but in the 2nd equation how did you getthat far..like ...break it waaay down 4 me =)

OpenStudy (anonymous):

2x-y=19---------------(i) 3y=2x-29-------------(ii) 2x-3y=29 (first step is to arrange the equations) 2x-y=19

OpenStudy (anonymous):

2x-3y=29 (subtract) -(2x-y)=-(19)

OpenStudy (anonymous):

2x-3y=29 -2x+y=-19 --------------------- 0-2y=-10

OpenStudy (anonymous):

-2y=10

OpenStudy (anonymous):

-y=10/2

OpenStudy (anonymous):

-y=5

OpenStudy (anonymous):

y=-5

OpenStudy (anonymous):

all clear?

OpenStudy (anonymous):

got it!